# Temperature dependence of coercivity for isolated Ni nanowires unraveled by high-sensitivity micromagnetometry

### v1

@@ IV Conclusions
We studied the hysteresis behavior of thin ferromagnetic Ni NWs with $(72\pm 5)$ nm average diameter and lengths ranging from $5$ to $9$ $\mu$m. After being released from the alumina template, eleven NWs were glued onto the plate of a micromechanical torsional oscillator, which was used as a micromagnetometer of high sensitivity. Since the interactions between these NWs are negligible, they can be considered to be in effect isolated. In this way, our measurements correspond to an average over this small group of non-interacting Ni NWs.
The hysteresis loops of $\Delta\nu$ vs $H$, obtained by applying the external field along the easy axis of the NWs, allows direct measurement of the coercivity. We found that $H_{c}$ does not change significantly over the temperature range from $5$ K to $200$ K, which contrasts strongly with the behavior previously observed for dense arrays of these same Ni NWs still embedded in the AAO template. This result confirms that the mechanical interaction between the nanowires with both, the template and the substrate, as well as the dipolar interactions between the NWs (which are absent in our experiments), are the main sources of distortion of coercivity. Also, when comparing with the Stoner-Wohlfarth model, it is clear that the magnetization of an isolated Ni NW is not reversed by coherent rotation.
@@ Acknowledgments
All raw data corresponding to the findings in this manuscript can be found in Ref. raw-data .
Appendix A
A silicon micromechanical torsional oscillator can be used as high sensitivity micromagnetometer. The resonant frequency of the microdevice with a magnetic sample adhered to its plate is
@@
where $k_{e}$ is the elastic restorative constant of the serpentine springs and $I$ is the moment of inertia of the whole system along its center rotational axis.
Because Ni NWs are ferromagnetic and have a large shape anisotropy, well below their Curie temperature an external magnetic field $\mathbf{H}$ applied parallel to the plate plane exerts an additional restoring torque and therefore the new resonant frequency will be
@@
Here, $k_{M}$ is the effective elastic constant originated by the interaction between the magnetization of the sample, $\mathbf{M}$, and the field, $\mathbf{H}$. As discussed in Sec. II.2, $k_{M}$ will be zero in cases I and II, positive in case III (restoring torque), and negative in case IV (anti-restoring torque).
To calculate the magnetization of the sample, we proceed as follows. Typically, $|\Delta\nu|=|\nu_{r}-\nu_{0}|\ll\nu_{0}$. Then, from Eqs.(5) and (6) it is possible to deduce that
@@
This effective elastic constant depends on $\mathbf{M}$ and it can be calculated from energy considerations Dolz2008 ; Zijlstra1961 ; Morillo1998 . Given the high aspect ratio of the NWs, the energy per unit volume of the sample can be written as
@@
where the first term represents the Zeeman interaction and the second is the uniaxial shape anisotropy energy. $\mathbf{n}$ is a unit vector pointing along the major (easy) axis of each NW (parallel to the plate of the micromechanical oscillator), and $K_{n}$ is the corresponding uniaxial shape anisotropy constant. As the oscillation’s amplitude is very small (around 1 sexagesimal degree at resonance), it is possible to write that Zijlstra1961 ; Morillo1998
@@
where $V_{n}$ is the volume of the sample and we have assumed that both, the temperature and the magnitude of the external field, are constants. Equation (9) is valid if the module of magnetization does not change appreciably throughout the oscillation cycle, a condition that is well satisfied given the small amplitude of oscillation.
Taking $K_{n}=\mu_{0}M^{2}/4$, the value of the shape anisotropy constant for an infinitely long rod Cullity , Eq. (9) can be written as a quadratic equation in $M$,
@@
which allows us to calculate the magnetization from the the experimental measurements of the change in resonance frequency, $\Delta\nu$. To guarantee that $dM/dH\geq 0$, the positive (negative) root of Eq. (12) should be taken when $\Delta\nu>0$ ($\Delta\nu<0$). This rule can be avoided by rewriting Eq. (12) as

### v2

@@ IV Conclusions
[…]
The hysteresis loops of $\Delta\nu$ vs $H$, obtained by applying the external field along the easy axis of the NWs, allows direct measurement of the coercivity. We found that $H_{c}$ does not change significantly over the temperature range from $5$ to $200$ K, which contrasts strongly with the behavior previously observed for dense arrays of these same Ni NWs still embedded in the AAO template. This result confirms that the mechanical interaction between the NWs with both the template and the substrate as well as the dipolar interactions between the NWs (which are absent in our experiments) is the main source of distortion of coercivity. Also, when comparing with the Stoner-Wohlfarth model, it is clear that the magnetization of an isolated Ni NW is not reversed by coherent rotation.
@@ Acknowledgments
The data that support the findings in this article are openly available raw-data .
Appendix A Magnetic hysteresis loops
A silicon micromechanical torsional oscillator can be used as a high-sensitivity micromagnetometer. The resonant frequency of the microdevice with a magnetic sample adhered to its plate is
@@ Appendix A Magnetic hysteresis loops
where $k_{e}$ is the elastic restorative constant of the serpentine springs, and II is the moment of inertia of the whole system along its center rotational axis.
Because Ni NWs are ferromagnetic and have a large shape anisotropy well below their Curie temperature, an external magnetic field $\mathbf{H}$ applied parallel to the plate plane exerts an additional restoring torque, and therefore, the new resonant frequency will be
@@
Here, $k_{M}$ is the effective elastic constant originated by the interaction between the magnetization of the sample $\mathbf{M}$ and the field $\mathbf{H}$. As discussed in Sec. II.2, $k_{M}$ will be zero in cases I and II, positive in case III (restoring torque), and negative in case IV (antirestoring torque).
To calculate the magnetization of the sample, we proceed as follows. Typically, $|\Delta\nu|=|\nu_{r}-\nu_{0}|\ll\nu_{0}$. Then from Eqs. (5) and (6), it is possible to deduce that
@@
This effective elastic constant depends on $\mathbf{M}$, and it can be calculated from energy considerations Dolz2008 ; Zijlstra1961 ; Morillo1998 . Given the high aspect ratio of the NWs, the energy per unit volume of the sample can be written as
@@
where the first term represents the Zeeman interaction, and the second is the uniaxial shape anisotropy energy. Here, $\mathbf{n}$ is a unit vector pointing along the major (easy) axis of each NW (parallel to the plate of the micromechanical oscillator), and $K_{n}$ is the corresponding uniaxial shape anisotropy constant. As the oscillation amplitude is very small ($\sim 1$ sexagesimal degree at resonance), it is possible to write that Zijlstra1961 ; Morillo1998
@@
where $V_{n}$ is the volume of the sample, and we have assumed that both the temperature and the magnitude of the external field are constants. Equation (9) is valid if the module of magnetization does not change appreciably throughout the oscillation cycle, a condition that is well satisfied given the small amplitude of oscillation.
Taking $K_{n}=\mu_{0}M^{2}/4$, the value of the shape anisotropy constant for an infinitely long rod Cullity , Eq. (9) can be written as a quadratic equation in $M$
@@
which allows us to calculate the magnetization from the the experimental measurements of the change in resonance frequency $\Delta\nu$. To guarantee that $dM/dH\geq 0$, the positive (negative) root of Eq. (12) should be taken when $\Delta\nu>0$ ($\Delta\nu<0$). This rule can be avoided by rewriting Eq. (12) as
@@
Appendix B Microdevice performance
The high sensitivity of the micromechanical oscillator is mainly due to the fact that the microdevice has a very high quality factor $Q$. We define this quantity as $Q=\nu_{0}/\Delta\nu_{0}$ where, as before, $\nu_{0}$ is the resonance frequency. Here, $\Delta\nu_{0}$ is the bandwidth (or resonance width) over which the vibration power (which is proportional to the square of the oscillation amplitude) is greater than half the power at the resonant frequency (do not mistake this quantity with $\Delta\nu$, which represents the change in resonant frequency due to the application of an external field and the subsequent change in the state of magnetization of the sample).
Figure 8: Dependence of the quality factor $Q$ and the resonance frequency $\nu_{0}$ (inset) with temperature.
Figure 8 shows the dependence of $Q$ and $\nu_{0}$ (inset) with temperature. While between $5$ and $300$ K the resonant frequency changes only $\sim 200$ Hz ($0.3\%$ of its value), the quality factor shows more pronounced changes: It reaches very high values at low temperatures but then decreases by an order of magnitude and stabilizes at $\sim 1.5\times 10^{4}$ at temperatures $\geq 125$ K.
Although the optimal performance of the microdevice evidently takes place in the low-temperature range, the values of $Q$ for $T\gtrsim 125$ K are still large enough to detect significant changes in the magnetization of the sample. However, in this work, measurements with adequate precision could only be made up to temperatures of $200$ K. The reason is that, although the sensitivity of the device depends on $Q$, it is also affected by thermal fluctuations.
Figure 9: Temperature dependence of the uncertainty values for different physical magnitudes. The frequency variation uncertainty $\Delta\nu_{\textrm{unc}}$ (open squares), obtained from a direct measurement, is constant for each temperature. In contrast, the magnetization uncertainty $\mu_{0}M_{\textrm{unc}}$, derived indirectly, spans a range between minimum (solid triangles) and maximum (solid circles) values.
To quantify the oscillator sensitivity, we first considered the uncertainty of a typical measurement of the change in the resonant frequency $\Delta\nu$ at each temperature $\Delta\nu_{\textrm{unc}}$. Figure 9 shows that this quantity increases exponentially with $T$ and reaches a value of $0.2$ Hz at room temperature (measurements made at $T\approx 300$ K, although not useful for accurately determining the coercive field, have been useful to estimate the sensitivity of the microdevice at this temperature). Since $\Delta\nu_{\textrm{unc}}\approx 0.1$ Hz at $T=200$ K, at first glance, it would seem that by averaging over a larger number of measurements, it would be possible to make a sufficiently accurate estimate of the magnetization at room temperature.
This last observation is incorrect since our magnetometer does not allow direct measurement of magnetization. Instead, for each measurement of $\Delta\nu$, it is necessary to use Eq. (2) to calculate a single value of $M$, and it is these magnetization values that must ultimately be averaged. It is worth noting that Eq. (2) includes the $\Delta\nu/H$ and $H^{2}/\Delta\nu$ ratios, which have a significant effect on $M$ when $\Delta\nu\to 0$, i.e., when the applied field $H\to 0$ (case I in Fig. 4) or when the coercive field is reached (case II).
We have therefore calculated $\mu_{0}M_{\textrm{unc}}$, the uncertainty resulting from this indirect measurement of the magnetization. Figure 9 also shows the temperature dependence of the maximum (which arises for cases I and II) and minimum (for cases III and IV) values that this quantity takes. As we can see, at room temperature, the uncertainty in the magnetization can reach $\mu_{0}M_{\textrm{unc}}\gtrsim 1$ T. Considering that the saturation magnetization is $\mu_{0}M_{s}=(0.6\pm 0.1)$ T, this result shows that it is not possible to obtain a complete magnetization loop at this temperature using our micromagnetometer.
Finally, note that between $5$ and $200$ K, $\mu_{0}M_{\textrm{unc}}$ ranges from $\sim 10^{-3}$ to $10^{-1}$ T. Considering the volume of the sample, these values suggest that the changes in magnetic moment that can be detected with our oscillator range from $10^{-15}$ to $10^{-13}$ A m2.